OpenStudy (anonymous):

Suppose you have a six sided die, with sides number 1 through 6, and you roll it several times.What is the probability that the first time you observe a 1 is on the second roll?

OpenStudy (anonymous):

whats ur choices??

OpenStudy (anonymous):

We are not given any other choices.The question is as I have mentioned

OpenStudy (anonymous):

2/6

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

How?

OpenStudy (raden):

how many times the die have rolled, u have to mention it

OpenStudy (anonymous):

That's what is confusing me. It's not mentiooned in the question.

OpenStudy (raden):

if just 2 times, it means the propability for the first roll, the event not 1 = 5/6 for the 2nd roll, the event 1 = 1/6 so, the event of it = 5/6 * 1/6 = 5/36

OpenStudy (anonymous):

And similarly if I am to calculate it for the third roll,it should be 5/6 * 5/6 * 1/6=25/216

OpenStudy (raden):

yes...

OpenStudy (raden):

because the events are independent, so just multiply of them

OpenStudy (anonymous):

Thank You!..

OpenStudy (raden):

welcome